# GLGENN: A Novel Parameter-Light Equivariant Neural Networks Architecture Based on Clifford Geometric Algebras

### v1

@@ F.4 Relation Between Ordinary and Generalized Lipschitz Groups
Therefore, by the definition (74) of the twisted centralizers, the elements $\widehat{\widetilde{T}}T$ and $\widetilde{T}T$ belong to the twisted centralizer of the grade-$1$ subspace $C\kern-1.00006pt\ell^{1}_{p,q,r}$ in $C\kern-1.00006pt\ell_{p,q,r}$:
@@ Appendix H Degenerate and Non-degenerate Lipschitz, Generalized Lipschitz, and Orthogonal Groups Equivariance
${\rm O}(V,\mathfrak{q})\cong\{\begin{pmatrix}A&0_{(p+q)\times r}\\ M&G\end{pmatrix}\},\qquad A\in{\rm O}(C\kern-1.00006pt\ell^{1}_{p,q},\mathfrak{q}|_{C\kern-0.70004pt\ell^{1}_{p,q}}),\qquad M\in{\rm Mat}_{r\times(p+q)}({\mathbb{F}}),\qquad G\in{\rm GL}(r,{\mathbb{F}});$ (190)
${\rm O}_{\Lambda^{1}_{r}}(V,\mathfrak{q})\cong\{\begin{pmatrix}A&0_{(p+q)\times r}\\ M&\mathrm{I}_{r}\end{pmatrix}\},\qquad A\in{\rm O}(C\kern-1.00006pt\ell^{1}_{p,q},\mathfrak{q}|_{C\kern-0.70004pt\ell^{1}_{p,q}}),\qquad M\in{\rm Mat}_{r\times(p+q)}({\mathbb{F}}),$ (191)
where ${\rm O}(C\kern-1.00006pt\ell^{1}_{p,q},\mathfrak{q}|_{C\kern-0.70004pt\ell^{1}_{p,q}})$ is the non-degenerate orthogonal group of transformations of $C\kern-1.00006pt\ell^{1}_{p,q}$ with a non-degenerate quadratic form $\mathfrak{q}|_{C\kern-0.70004pt\ell^{1}_{p,q}}$; $0_{(p+q)\times r}$ is the zero matrix of size $(p+q)\times r$; ${\rm Mat}_{r\times(p+q)}({\mathbb{F}})$ is a set of arbitrary matrices of size $r\times(p+q)$ with coefficients in ${\mathbb{F}}$; $\mathrm{I}_{r}$ is the identity matrix of size $r\times r$; ${\rm GL}(r,{\mathbb{F}})$ is the general linear group acting on an $r$-dimensional vector space over ${\mathbb{F}}$.
@@
$\ker(\tilde{{\rm ad}}^{1}:\tilde{\Gamma}^{1}_{p,q,r}\rightarrow{\rm Aut}(C\kern-1.00006pt\ell_{p,q,r}))$ =$=\{T\in\tilde{\Gamma}^{1}_{p,q,r}:\quad\widehat{T}vT^{-1}=v,\quad\forall v\in C\kern-1.00006pt\ell^{1}_{p,q,r}\}$ (193)
@@
Let us prove these statement. Firstly, $U$ is invertible, since $(e+me_{ab})(e-me_{ab})=e$. Secondly $U$ preserves the grade-$1$ subspace under $\tilde{{\rm ad}}$, since
@@
where $1$ is on the $k$-th position. On the other hand, for $U_{1,1}$ and $e_{k}$, we get
@@
where we apply Remark H.3, since $\tilde{{\rm ad}}_{T}(v)\in V$, and multiplicativity of $\tilde{{\rm ad}}_{T}$ (Lemma 2.1), since $T\in\tilde{\Gamma}^{1}_{p,q,r}\subseteq(C\kern-1.00006pt\ell^{(0)\times}_{p,q,r}\cup C\kern-1.00006pt\ell^{(1)\times}_{p,q,r})\Lambda^{\times}_{r}$ by Theorem 3.3. Since $v^{2}=\mathfrak{q}(v)$ again by Remark H.3, we finally get
@@
${\rm O}_{\Lambda^{1}_{r}}(V,\mathfrak{q})\cong\{\begin{pmatrix}A&0\\ M&\mathrm{I}_{r}\end{pmatrix}\},\qquad A\in{\rm O}(C\kern-1.00006pt\ell^{1}_{p,q},\mathfrak{q}|_{C\kern-0.70004pt\ell^{1}_{p,q}}),\qquad M\in{\rm Mat}_{r\times(p+q)}({\mathbb{F}}).$ (217)
Suppose $\Phi\in{\rm O}_{\Lambda^{1}_{r}}(V,\mathfrak{q})$. Then, there exist $A\in{\rm O}(C\kern-1.00006pt\ell^{1}_{p,q},\mathfrak{q}|_{C\kern-0.70004pt\ell^{1}_{p,q}})$ and $M\in{\rm Mat}_{r\times(p+q)}({\mathbb{F}})$ such that
@@
where $A=A_{1}\cdots A_{k}$, $k\leq p+q$, and $A_{1},\ldots,A_{k}\in{\rm O}(C\kern-1.00006pt\ell^{1}_{p,q},\mathfrak{q}|_{C\kern-0.70004pt\ell^{1}_{p,q}})$ are matrices representing reflections, by the well-known Cartan–Dieudonné theorem (Dieudonné, 1971), which states that every orthogonal transformation in a $(p+q)$-dimensional space with a non-degenerate symmetric bilinear form can be represented as a composition of at most $p+q$ reflections; and $M_{i,j}\in{\rm Mat}_{r\times(p+q)}({\mathbb{F}})$, $i=1,\ldots,r$, $j=1,\ldots,p+q$ is a matrix with at most one non-zero element, which is on the $i$-th row and $j$-th column, such that $M=\sum_{i=1,\ldots,r}\sum_{j=1,\ldots,p+q}M_{i,j}$.
@@
[…]
@@ Appendix I Experimental Details
We construct our models to closely resemble the CGENN architecture, replacing the $C\kern-1.00006pt\ell^{k}_{p,q}$-linear, $C\kern-1.00006pt\ell^{k}_{p,q}$-geometric product, and $C\kern-1.00006pt\ell^{k}_{p,q}$-normalization layers from Ruhe et al., 2023 with the same number of GLGENN’s $C\kern-1.00006pt\ell^{\overline{k}}_{p,q}$ counterparts (see Section 4). This design choice results in GLGENN having significantly fewer parameters than CGENN, as they operate in a unified manner across $4$ fundamental subspaces of geometric algebras defined by the grade involution ($\widehat{\;\;}$) and reversion ($\widetilde{\;\;}$); they processes geometric objects in groups with a step size of $4$.
However, this parameter efficiency difference between GLGENN and CGENN becomes apparent only for tasks with $n>3$. In lower-dimensional cases ($n\leq 3$), subspaces of fixed grades coincide with the subspaces determined by grade involution and reversion (4), i.e., $C\kern-1.00006pt\ell^{k}_{p,q,r}=C\kern-1.00006pt\ell^{\overline{k}}_{p,q,r}$ for $k=0,1,2,3$. Nonetheless, as experiments show, even in $5$-dimensional cases, the performance gap between CGENN and GLGENN is substantial. Moreover, GLGENN’s parameter efficiency advantage increases with dimension $n$, as the subspaces $C\kern-1.00006pt\ell^{\overline{k}}_{p,q,r}$ diverge further from $C\kern-1.00006pt\ell^{k}_{p,q,r}$.
@@
In the case $K=16$, we consider $4$ different sizes of the training set: $256$, $1024$, $4096$, and $16384$ samples. We compare GLGENN with the state-of-the-art model CGENN (Ruhe et al., 2023). While both models share a similar architecture, GLGENN has 24.1K trainable parameters, whereas CGENN has 58.8K. The number of training steps is the same for both models. The results are presented in Table 3 and Figure 2 (middle). GLGENN either outperforms or matches CGENN across all training set sizes. The most noticeable difference occurs with smaller training sets. We attribute this to GLGENN’s lower parameter count, which reduces its tendency to overfit—an issue commonly observed in small datasets.
@@
GLGENN $0.6$K 0.1055 0.0020 0.0031 0.0011
CGENN (same size as GLGENN) $0.6$K 0.2856 0.0076 0.0017 0.0005
CGENN (best performing) $1.8$K 0.0791 0.0089 0.0012 0.0003
@@
Table 12: MSE ($\downarrow$) on the ${\rm O}(5,0)$-$N$-Body Experiment.

### v2

@@ F.4 Relation Between Ordinary and Generalized Lipschitz Groups
Therefore, by the definition (74) of the twisted centralizers, the elements $\widehat{\widetilde{T}}T$ and $\widetilde{T}T$ belong to the twisted centralizer of the grade-11 subspace $C\kern-1.00006pt\ell^{1}_{p,q,r}$ in $C\kern-1.00006pt\ell_{p,q,r}$:
@@ Appendix H Degenerate and Non-degenerate Lipschitz, Generalized Lipschitz, and Orthogonal Groups Equivariance
${\rm O}(V,\mathfrak{q})\cong\{\begin{pmatrix}A&0_{(p+q)\times r}\\ M&G\end{pmatrix}\},\qquad A\in{\rm O}(C\kern-1.00006pt\ell^{1}_{p,q},\mathfrak{q}|_{C\kern-0.81949pt\ell^{1}_{p,q}}),\qquad M\in{\rm Mat}_{r\times(p+q)}({\mathbb{F}}),\qquad G\in{\rm GL}(r,{\mathbb{F}});$ (190)
${\rm O}_{\Lambda^{1}_{r}}(V,\mathfrak{q})\cong\{\begin{pmatrix}A&0_{(p+q)\times r}\\ M&\mathrm{I}_{r}\end{pmatrix}\},\qquad A\in{\rm O}(C\kern-1.00006pt\ell^{1}_{p,q},\mathfrak{q}|_{C\kern-0.81949pt\ell^{1}_{p,q}}),\qquad M\in{\rm Mat}_{r\times(p+q)}({\mathbb{F}}),$ (191)
where ${\rm O}(C\kern-1.00006pt\ell^{1}_{p,q},\mathfrak{q}|_{C\kern-0.81949pt\ell^{1}_{p,q}})$ is the non-degenerate orthogonal group of transformations of $C\kern-1.00006pt\ell^{1}_{p,q}$ with a non-degenerate quadratic form $\mathfrak{q}|_{C\kern-0.81949pt\ell^{1}_{p,q}}$; $0_{(p+q)\times r}$ is the zero matrix of size $(p+q)\times r$; ${\rm Mat}_{r\times(p+q)}({\mathbb{F}})$ is a set of arbitrary matrices of size $r\times(p+q)$ with coefficients in $\mathbb{F}$}; $\mathrm{I}_{r}$ is the identity matrix of size $r\times r$; ${\rm GL}(r,{\mathbb{F}})$ is the general linear group acting on an $r$-dimensional vector space over $\mathbb{F}$}.
@@
$\ker(\tilde{{\rm ad}}^{1}:\tilde{\Gamma}^{1}_{p,q,r}\rightarrow{\rm Aut}(C\kern-1.00006pt\ell_{p,q,r}))$ $=\{T\in\tilde{\Gamma}^{1}_{p,q,r}:\quad\widehat{T}vT^{-1}=v,\quad\forall v\in C\kern-1.00006pt\ell^{1}_{p,q,r}\}$ (193)
@@
Let us prove these statement. Firstly, $U$ is invertible, since $(e+me_{ab})(e-me_{ab})=e$. Secondly $U$ preserves the grade-11 subspace under $\tilde{{\rm ad}}$, since
@@
where 11 is on the $k$-th position. On the other hand, for $U_{1,1}$ and $e_{k}$, we get
@@
where we apply Remark H.3, since $\tilde{{\rm ad}}_{T}(v)\in V$, and multiplicativity of $\tilde{{\rm ad}}_{T}$ (Lemma 2.1), since $T\in\tilde{\Gamma}^{1}_{p,q,r}\subseteq(C\kern-1.00006pt\ell^{(0)\times}_{p,q,r}\cup C\kern-1.00006pt\ell^{(1)\times}_{p,q,r})\Lambda^{\times}_{r}$ by Theorem $3.$ Since $v^{2}=\mathfrak{q}(v)$ again by Remark H.3, we finally get
@@
${\rm O}_{\Lambda^{1}_{r}}(V,\mathfrak{q})\cong\{\begin{pmatrix}A&0\\ M&\mathrm{I}_{r}\end{pmatrix}\},\qquad A\in{\rm O}(C\kern-1.00006pt\ell^{1}_{p,q},\mathfrak{q}|_{C\kern-0.81949pt\ell^{1}_{p,q}}),\qquad M\in{\rm Mat}_{r\times(p+q)}({\mathbb{F}}).$ (217)
Suppose $\Phi\in{\rm O}_{\Lambda^{1}_{r}}(V,\mathfrak{q})$. Then, there exist $A\in{\rm O}(C\kern-1.00006pt\ell^{1}_{p,q},\mathfrak{q}|_{C\kern-0.81949pt\ell^{1}_{p,q}})$ and $M\in{\rm Mat}_{r\times(p+q)}({\mathbb{F}})$ such that
@@
where $A=A_{1}\cdots A_{k}$, $k\leq p+q$, and $A_{1},\ldots,A_{k}\in{\rm O}(C\kern-1.00006pt\ell^{1}_{p,q},\mathfrak{q}|_{C\kern-0.81949pt\ell^{1}_{p,q}})$ are matrices representing reflections, by the well-known Cartan–Dieudonné theorem (Dieudonné, 1971), which states that every orthogonal transformation in a $(p+q)$-dimensional space with a non-degenerate symmetric bilinear form can be represented as a composition of at most $p+q$ reflections; and $M_{i,j}\in{\rm Mat}_{r\times(p+q)}({\mathbb{F}})$, $i=1,\ldots,r$, $j=1,\ldots,p+q$ is a matrix with at most one non-zero element, which is on the $i$-th row and $j$-th column, such that $M=\sum_{i=1,\ldots,r}\sum_{j=1,\ldots,p+q}M_{i,j}$.
@@
$\tilde{{\rm ad}}^{1}:\quad{\raisebox{2.04439pt}{$\tilde{\Gamma}^{1}_{p,q,r}$}\left/\raisebox{-2.04439pt}{$\Lambda_{r}^{\times}$}\right.}\cong{\rm O}_{\Lambda^{1}_{r}}(V,\mathfrak{q}),$ (227)
@@
We have the homomorphism $\tilde{{\rm ad}}^{1}$ of the groups $\tilde{\Gamma}^{1}_{p,q,r}$ and ${\rm O}_{\Lambda^{1}_{r}}(V,\mathfrak{q})$, which is surjective by the statement (212) of Theorem H.9. By the fundamental homomorphism theorem, the group ${\rm O}_{\Lambda^{1}_{r}}(V,\mathfrak{q})$ is isomorphic to the quotient group $\raisebox{1.99997pt}{$\tilde{\Gamma}^{1}_{p,q,r}$}\left/\raisebox{-1.99997pt}{$\ker(\tilde{{\rm ad}}^{1})$}\right$.}. By applying Theorem H.2, we complete the proof. ∎
@@ Appendix I Experimental Details
We construct our models to closely resemble the CGENN architecture, replacing the $C\kern-1.00006pt\ell^{k}_{p,q}$-linear, $C\kern-1.00006pt\ell^{k}_{p,q}$-geometric product, and $C\kern-1.00006pt\ell^{k}_{p,q}$-normalization layers from Ruhe et al., 2023 with the same number of GLGENN’s $C\kern-1.00006pt\ell^{\overline{k}}_{p,q}$ counterparts (see Section 4). This design choice results in GLGENN having significantly fewer parameters than CGENN, as they operate in a unified manner across 44 fundamental subspaces of geometric algebras defined by the grade involution ($\widehat{\;\;}$) and reversion ($\widetilde{\;\;}$); they processes geometric objects in groups with a step size of 44.
However, this parameter efficiency difference between GLGENN and CGENN becomes apparent only for tasks with $n>3$. In lower-dimensional cases ($n\leq 3$), subspaces of fixed grades coincide with the subspaces determined by grade involution and reversion (4), i.e., $C\kern-1.00006pt\ell^{k}_{p,q,r}=C\kern-1.00006pt\ell^{\overline{k}}_{p,q,r}$ for $k=0,1,2,3$. Nonetheless, as experiments show, even in 55-dimensional cases, the performance gap between CGENN and GLGENN is substantial. Moreover, GLGENN’s parameter efficiency advantage increases with dimension $n$, as the subspaces $C\kern-1.00006pt\ell^{\overline{k}}_{p,q,r}$ diverge further from $C\kern-1.00006pt\ell^{k}_{p,q,r}$.
@@
In the case $K=16$, we consider 44 different sizes of the training set: $256$, $1024$, $4096$, and $16384$ samples. We compare GLGENN with the state-of-the-art model CGENN (Ruhe et al., 2023). While both models share a similar architecture, GLGENN has 24.1K trainable parameters, whereas CGENN has 58.8K. The number of training steps is the same for both models. The results are presented in Table 3 and Figure 2 (middle). GLGENN either outperforms or matches CGENN across all training set sizes. The most noticeable difference occurs with smaller training sets. We attribute this to GLGENN’s lower parameter count, which reduces its tendency to overfit—an issue commonly observed in small datasets.
@@
GLGENN 0.60.6K 0.1055 0.0020 0.0031 0.0011
CGENN (same size as GLGENN) 0.60.6K 0.2856 0.0076 0.0017 0.0005
CGENN (best performing) 1.81.8K 0.0791 0.0089 0.0012 0.0003
@@
Table 12: MSE ($\downarrow$) on the ${\rm O}(5,0)$-NN-Body Experiment.